# Text-controlled Motion Mamba: Text-Instructed Temporal Grounding of Human Motion

### v1

@@ 2.3. State Space Model
3:$\mathbf{B}$ : $(\mathtt{V},\mathtt{L},\mathtt{N})\leftarrow\mathrm{Linear}_{\mathbf{B}}(\mathbf{X})$
4:$\mathbf{C}$ : $(\mathtt{V},\mathtt{L},\mathtt{N})\leftarrow\mathrm{Linear}_{\mathbf{C}}(\mathbf{X})$
5:$\mathbf{\Delta}$ : $(\mathtt{V},\mathtt{L},\mathtt{D})\leftarrow\log(1+\exp(\mathrm{Linear}_{\mathbf{\Delta}}(\mathbf{X})+\mathrm{Parameter}_{\mathbf{\Delta}}))$
6:$\overline{\mathbf{A}},\overline{\mathbf{B}}$ : $(\mathtt{V},\mathtt{L},\mathtt{D},\mathtt{N})\leftarrow\mathrm{discretize}(\mathbf{\Delta},\mathrm{Parameter}_{\mathbf{A}},\mathbf{B})$
7:$\mathbf{Y}\leftarrow\mathrm{SSM}(\overline{\mathbf{A}},\overline{\mathbf{B}},\mathbf{C})(\mathbf{X})$
@@
1:Input sequence $\mathbf{X}\in\mathbb{R}^{\mathtt{V}\times\mathtt{L}\times\mathtt{D}}$; Text query embedding $q\in\mathbb{R}^{D}$
@@
3:$\mathbf{B}$ : $(\mathtt{V},\mathtt{L},\mathtt{N})\leftarrow\mathrm{Linear}_{\mathbf{B}}(\mathbf{X},{\color[rgb]{0.0,0.5,0.0}q})$
4:$\mathbf{C}$ : $(\mathtt{V},\mathtt{L},\mathtt{N})\leftarrow\mathrm{Linear}_{\mathbf{C}}(\mathbf{X},{\color[rgb]{0.0,0.5,0.0}q})$
5:$\mathbf{\Delta}$ : $(\mathtt{V},\mathtt{L},\mathtt{D})\leftarrow\log(1+\exp(\mathrm{Linear}_{\mathbf{\Delta}}(\mathbf{X},{\color[rgb]{0.0,0.5,0.0}q})+\mathrm{Parameter}_{\mathbf{\Delta}}))$
6:$\overline{\mathbf{A}},\overline{\mathbf{B}}$ : $(\mathtt{V},\mathtt{L},\mathtt{D},\mathtt{N})\leftarrow\mathrm{discretize}(\mathbf{\Delta},\mathrm{Parameter}_{\mathbf{A}},\mathbf{B})$
7:$\mathbf{Y}\leftarrow\mathrm{SSM}(\overline{\mathbf{A}},\overline{\mathbf{B}},\mathbf{C})(\mathbf{X})$
@@
3. BABEL-Grounding Dataset
The task of THMG involves determining the start and end timestamps of all the segments in the motion sequence that align with the given textual description. Our objective is to construct a dataset where each text query depicting specific human actions can be mapped to one or more temporal segments within the motion sequence. To this end, we construct the BABEL-Grounding dataset based on BABEL (Punnakkal et al., 2021). BABEL-Grounding provides detailed textual descriptions of human actions alongside their corresponding temporal segments in untrimmed motion sequences, with a total of 5,339 sequences with 21,307 text-segments annotations. Each sequence averages 743 frames, with the ground-truth temporal segments having an average frame count of 112, indicating their sparse distribution within the lengthy motion sequences. Figure 2 illustrates the dataset statistics of BABEL-Grounding. As depicted, the BABEL-Grounding dataset contains comprehensive textual descriptions for motion sequences of diverse lengths. Each text query may correspond to multiple temporal segments, and these segments are sparsely distributed throughout the entire motion sequence, inherently posing a challenge for motion grounding task.
Below, we provide a concise overview of the construction process for BABEL-Grounding annotations. The original BABEL dataset provides dense temporal annotations by labeling each temporal segment with corresponding actions. However, its textual annotations consist of simple categorical phrases rather than detailed and comprehensive sentences, and its data structure is not directly applicable to the THMG task. Therefore, a set of manually-crafted rules has been employed to augment the dataset, which will be elaborated below.
3.1. Textual augmentation
Utilizing external annotations.
BABEL is built upon the AMASS motion capture database (Mahmood et al., 2019), while HumanML3D (Guo et al., 2022a) offers detailed textual annotations at the sequence level for the AMASS database. For each entry in the BABEL dataset with overly simplistic annotations, its corresponding entry in HumanML3D is located by the sequence ID. When the sequence-level HumanML3D annotations contain the phrase from the BABEL annotations, human annotators will manually verify their correspondence and supplement the BABEL annotations with detailed textual descriptions from HumanML3D.
Template-based augmentation.
As HumanML3D’s sequence-level annotations only partially cover the items in BABEL, many low-quality annotations remain unaddressed, particularly those with only one-word labels. To address this gap, we’ve manually crafted templates to enhance them. For example, ‘take’ is expanded to ‘take something with his hands’, ‘stir’ becomes ‘stir something with his hands in a circular motion’, and ‘place’ becomes ‘place the object at a specific location’. This substitution process enriches the overly simplistic annotations, rendering them more comprehensible for the model. The final textual annotations are further refined by ChatGPT(gpt-3.5-turbo) to make them more fluent, complete and diverse.
3.2. Temporal Augmentation
Time windows merging.
one-to-many mapping.
Figure 3. Left: overall architecture of our proposed model. Right: Illustration of TM-Mamba block. ‘Bidirectional SSM’ refers to the text-controlled selection mechanism demonstrated in Algorithm 2 with bidirectional modeling.
4. Method
4.1. Preliminaries on Mamba
In this section, we present a brief review of State Space Models (SSM) and Mamba (Gu and Dao, 2023). SSM is a series of sequential models that maps the input sequence $x(t)\in\mathbb{R}$ to the output sequence $y(t)\in\mathbb{R}$ through a hidden state $h(t)\in\mathbb{R}^{N}$, which can be depicted as a linear ODE:
(1) $h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t),$
@@ 4.1. Preliminaries on Mamba
(2) $\mathbf{\overline{A}}$ $=\exp{(\mathbf{\Delta}\mathbf{A})},$
@@
(3) $h_{t}$ $=\mathbf{\overline{A}}h_{t-1}+\mathbf{\overline{B}}x_{t},$
@@
However, linear recurrence requires unfolding in time and is unable to be parallelized. S4 (Gu et al., 2021) ensures Linear Time Invariance (LTI) by assuming that $\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{\Delta}$ remain static, allowing for the implementation using global convolution as $\mathbf{y}=\mathbf{x}*\mathbf{\overline{K}}$, where
(4) $\mathbf{\overline{K}}$ $=(\mathbf{C}\mathbf{\overline{B}},\mathbf{C}\mathbf{\overline{A}}\mathbf{\overline{B}},\dots,\mathbf{C}\mathbf{\overline{A}}^{L-1}\mathbf{\overline{B}}),$
and $L$ denotes the length of the input sequence, and $\overline{\mathbf{K}}\in\mathbb{R}^{L}$ represents a structured convolutional kernel. On the other hand, Mamba (Gu and Dao, 2023) introduces an input-dependent selection mechanism by making $\mathbf{\overline{A}},\mathbf{\overline{B}},\mathbf{C},\mathbf{\Delta}$ become functions of $x_{t}$. Such formulation can be efficiently computed via the proposed parallel scan algorithm.
4.2. Text-Controlled Selection Mechanism
@@
However, in the context of the THMG task, the model needs to dynamically select relevant global information from the sequence based on the text query to achieve better grounding performance. Existing multimodal Mamba methods (Qiao et al., 2024; Zhao et al., 2024) use simple concatenation to merge textual and visual features as input for Mamba, but the key selection mechanism governing the information flow remains the same. To overcome this limitation, we propose a text-controlled selection mechanism that allows the selection process to depend on both the motion input and the text query. As illustrated in Algorithm 1 and 2, Mamba parameterizes $\mathbf{\overline{A}},\mathbf{\overline{B}},\mathbf{C},\mathbf{\Delta}$ as functions of the input, whereas in text-controlled selection, these parameters become functions of input sequence as well as text query.
The algorithm essentially resembles a text-based gating mechanism, dynamically controlling the flow of information based on textual queries. Theorem 1 of (Gu and Dao, 2023) implies that Algorithm 2 exhibits similarities to gated RNN under certain conditions:
Lemma 4.1.
@@
Lemma 4.1 indicates that the text-controlled selective SSM bears resemblance to a gated RNN, wherein the gate $g_{k}$ relies on both motion input and text query, enabling the text query to control the information flow during propagation. Given that the recurrence process in line 5 of Algorithm 2 remains unchanged, the resultant text-controlled SSM can still be efficiently computed using the parallel scan algorithm outlined in (Gu and Dao, 2023). The new algorithm can be implemented by modifying the forward and gradient backward functions of the original Mamba, which enables end-to-end joint training of SSM and the language backbone.
Table 1. Ablation studies on BABEL-Grounding dataset. The best results are in bold.
@@
4.3. Text-Controlled Motion Mamba
@@
In this study, we enhance Mamba through the incorporation of topology awareness, achieved by integrating relational embeddings to convey information regarding neighboring nodes. Suppose an input sequence takes the shape of $\mathbf{X}\in\mathbb{R}^{V\times L\times D}$, where $V$ denotes the number of joints in the human skeleton, $L$ denotes the sequence length, and $D$ denotes the feature dimension. The relational embedding $\mathbf{R}$ is computed as $\mathbf{R}=f(\mathbf{X})\in\mathbb{R}^{V\times L\times D}$, which encapsulates the graph message at each time step. Here, $f$ denotes a Graph Neural Network (GNN), implemented as the AGCN proposed in  (Shi et al., 2019). Subsequently, the relational embedding $\mathbf{R}$ is concatenated with the motion features and fed into the text-controlled selective SSM in Algorithm 2, to jointly capture the global temporal information of each joint alongside its topological context.
The overall architecture of the Text-Controlled Motion Mamba is depicted in Figure 3. Unlike the vanilla Mamba block, which adopts unidirectional causal modeling, the task of THMG demands the global context of the entire sequence. To address this issue, a bidirectional non-causal structure, as proposed in Vision Mamba (Zhu et al., 2024), is employed. The input goes through a stack of TM-Mamba blocks, producing the output $\mathbf{Y}\in\mathbb{R}^{V\times L\times D}$. After mean pooling along the $V$ dimension, $\mathbf{Y}$ is subsequently forwarded to an MLP layer. This yields the frame activation score $s_{t}(t=1,2,\ldots,T)$ for each frame, indicating the likelihood of its inclusion in the retrieved temporal segments. The entire framework can be supervised using a simple cross-entropy loss.
(5) $\mathcal{L}_{ce}=-\frac{1}{T}\sum_{t}^{T}(y_{t}\log{s_{t}}+(1-y_{t})\log{(1-s_{t})}).$
@@
5. Experiments
5.1. Implementation and Evaluation Details
The motion data pre-processing procedure adheres to the methodology outlined in BABEL (Punnakkal et al., 2021). The maximum length of the motion sequence is constrained to 2000. The same data split for training and evaluation as BABEL is adopted. We employ CLIP (Radford et al., 2021) (clip-vit-base-patch32) as the text encoder, with its parameters being jointly tuned together with the entire model during training. The feature dimension $D$ in Algorithm 2 for both the input motion sequence and text query embedding is set to 256. The batch size and base learning rate are configured as $4$ and $5\times 10^{-4}$ respectively, with the learning rate for CLIP set to $5\times 10^{-5}$. The optimizer is an AdamW (Loshchilov and Hutter, 2017) with a weight decay of $1\times 10^{-4}$. The number of stacked Motion-Mamba blocks is empirically specified as $3$. Model training is conducted on a single Nvidia A40 GPU.
During the inference stage, a series of thresholds are used to obtain the predicted temporal segments, following (Caba Heilbron et al., 2015). Subsequently, non-maximum suppression is performed to remove overlapping segments. The evaluation of grounding performance is conducted using mean Average Precisions (mAPs) under different Intersection of Union (IoU) thresholds, namely $[0.1:0.7:0.1]$.
5.2. Ablation Studies
@@
Text-Controlled Selection Mechanism. The text-controlled selection mechanism lies at the core of the TM-Mamba model. Removing this component directly from Algorithm 2 for ablation causes the model incapable of perceiving text query information. To enable comparison, the ablative models follow the practice of (Qiao et al., 2024; Zhao et al., 2024), where textual embeddings are concatenated with the sequence input. An MLP is then employed for feature fusion. As demonstrated in Table 1, the text-controlled selection mechanism accounts for a substantial performance gain. This underscores the significance of dynamically regulating the propagation of information based on textual queries.
Unidirectional v.s. Bidirectional. The Vanilla Mamba model adopts a unidirectional approach, whereby the model can only access the sequence history while processing the current step. This unidirectional approach proves advantageous for tasks such as language modeling. However, in the context of the THMG task, the model requires a comprehensive understanding of the global context of the entire sequence, demanding a non-causal bidirectional structure. The results presented in Table 1 substantiate this claim, demonstrating that the bidirectional model significantly outperforms its unidirectional counterpart.
@@
Table 2. Performance comparisons to baseline methods on BABEL-Grounding dataset. The best results are in bold.
@@
S4D-LegS (Gu et al., 2022a) 29.1 26.2 23.0 19.6 15.9 12.0 8.9 19.2
S4D-Lin (Gu et al., 2022a) 30.1 26.9 23.8 20.5 16.6 12.8 9.5 20.0
2s-AGCN (Shi et al., 2019) 31.2 24.7 20.5 17.1 14.0 10.5 7.8 18.0
InfoGCN (Chi et al., 2022) 49.5 42.5 36.1 30.5 26.0 20.8 15.3 31.5
MomentDETR (Lei et al., 2021) 51.1 46.0 39.1 32.9 26.8 20.5 13.6 32.9
EaTR (Jang et al., 2023) 53.4 48.5 43.6 37.3 31.0 23.5 16.2 36.2
STCAT (Jin et al., 2022) 47.1 44.2 40.7 37.1 33.3 28.9 23.6 36.4
@@
Figure 4. Comparison of memory consumption for TM-Mamba and its transformer and LSTM counterparts under varying motion sequence lengths. The transformer model runs out of GPU memory when the sequence length reaches 1200.
5.3. Comparison with Baselines
Given that the THMG task is novel and lacks existing works for comparison, several baseline methods are implemented for analysis. To begin, several prominent models are chosen from the video moment retrieval task: MomentDETR (Lei et al., 2021), EaTR (Jang et al., 2023) and STCAT (Jin et al., 2022). However, these models rely on global temporal transformers, which require quadratic computational memory and are impractical for handling the lengthy sequences in THMG task. As a remedy, their global temporal transformers are substituted with efficient GRUs which require only linear memory. We also implement two baselines based on powerful human motion backbones using spatial-temporal graph convolutions, namely 2s-AGCN (Shi et al., 2019) and InfoGCN (Chi et al., 2022). Text embedding is incorporated through concatenation and MLP fusion to integrate query information for text-based grounding. Finally, we implemented several baselines based on recent SSM-based models (Gu et al., 2022a) to demonstrate the advantages of TM-Mamba over other SSM methods. As illustrated in Table 2, TM-Mamba outperforms the baseline models in terms of mAP at various IoU thresholds, thereby demonstrating its effectiveness on the BABEL-Grounding.
Table 3. Performance comparison of TM-Mamba and its temporal transformer counterpart under different maximum sequence length.
@@ 5.3. Comparison with Baselines
Figure 5. Visualizations of predicted frame activation score. The solid red line denotes the predicted score of our full model, while the dashed blue and gray lines denote the model without text control and relational embeddings, respectively. The gray bars below illustrate the ground-truth temporal segments. Best viewed in color.
Figure 6. Visualizations of the grounding results of different models. ‘GT’ denotes ground-truth temporal segments corresponding to the text query. Our TM-Mamba demonstrates superior performance in motion grounding, in terms of the number of retrieved segments and the temporal boundaries of each segment. Best viewed in color.
5.4. More Analysis

### v2

@@ III-C2 Manual Evaluation
[…]
@@
1:Input sequence $\mathbf{X}\in\mathbb{R}^{\mathtt{V}\times\mathtt{L}\times\mathtt{D}}$; Text embedding $q\in\mathbb{R}^{D}$
@@
[…]
@@
IV Method
IV-A Preliminaries on Mamba
In this section, we present a brief review of State Space Models (SSM) and Mamba [57]. SSM is a series of sequential models that maps the input sequence $x(t)\in\mathbb{R}$ to the output sequence $y(t)\in\mathbb{R}$ through a hidden state $h(t)\in\mathbb{R}^{N}$, which can be depicted as a linear ODE:
$h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t),$ (1)
@@ IV-A Preliminaries on Mamba
$\mathbf{\overline{A}}$ $=\exp{(\mathbf{\Delta}\mathbf{A})},$ (2)
@@
$h_{t}$ $=\mathbf{\overline{A}}h_{t-1}+\mathbf{\overline{B}}x_{t},$ (3)
@@
However, linear recurrence requires unfolding in time and is unable to be parallelized. S4 [91] ensures Linear Time Invariance (LTI) by assuming that $\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{\Delta}$ remain static, allowing for the implementation using global convolution as $\mathbf{y}=\mathbf{x}*\mathbf{\overline{K}}$, where
$\mathbf{\overline{K}}$ $=(\mathbf{C}\mathbf{\overline{B}},\mathbf{C}\mathbf{\overline{A}}\mathbf{\overline{B}},\dots,\mathbf{C}\mathbf{\overline{A}}^{L-1}\mathbf{\overline{B}}),$ (4)
and $L$ denotes the length of the input sequence, and $\overline{\mathbf{K}}\in\mathbb{R}^{L}$ represents a structured convolutional kernel. On the other hand, Mamba [57] introduces an input-dependent selection mechanism by making $\mathbf{\overline{A}},\mathbf{\overline{B}},\mathbf{C},\mathbf{\Delta}$ become functions of $x_{t}$. Such formulation can be efficiently computed via the proposed parallel scan algorithm.
IV-B Text-Controlled Selection Mechanism
@@
However, in the context of the THMG task, the model needs to dynamically select relevant global information from the sequence based on the text query to achieve better grounding performance. Existing multimodal Mamba methods [58, 59] use simple concatenation to merge textual and visual features as input for Mamba, but the key selection mechanism governing the information flow remains the same. To overcome this limitation, we propose a text-controlled selection mechanism that allows the selection process to depend on both the motion input and the text query. Given the input text query, we first utilize CLIP [118] to acquire the text embedding $q$. As illustrated in Algorithm 1 and 2, Mamba parameterizes $\mathbf{\overline{A}},\mathbf{\overline{B}},\mathbf{C},\mathbf{\Delta}$ as functions of the input, whereas in text-controlled selection, these parameters become functions of input sequence as well as text embedding $q$.
The algorithm essentially resembles a text-based gating mechanism, dynamically controlling the flow of information based on textual queries. Theorem 1 of [57] implies that Algorithm 2 exhibits similarities to gated RNN under certain conditions:
Lemma 1.
@@
Lemma 1 indicates that the text-controlled selective SSM bears resemblance to a gated RNN, wherein the gate $g_{k}$ relies on both motion input and text query, enabling the text query to control the information flow during propagation. Given that the recurrence process in line 5 of Algorithm 2 remains unchanged, the resultant text-controlled SSM can still be efficiently computed using the parallel scan algorithm outlined in [57]. The new algorithm can be implemented by modifying the forward and gradient backward functions of the original Mamba, which enables end-to-end joint training of SSM and the language backbone.
TABLE II: Ablation studies on BABEL-Grounding dataset. The best results are in bold.
@@
IV-C Text-Controlled Motion Mamba
@@
In this study, we enhance Mamba through the incorporation of topology awareness, achieved by integrating relational embeddings to convey information regarding neighboring nodes. The input skeleton sequence is first embedded by a linear layer, resulting in an embedding sequence with shape $\mathbf{X}\in\mathbb{R}^{V\times L\times D}$, where $V$ denotes the number of joints in the human skeleton, $L$ denotes the sequence length, and $D$ denotes the feature dimension. The graph-structured representation of human skeleton has been extensively validated in prior research, with Graph Convolutional Networks (GCNs) demonstrating exceptional capability in modeling inter-joint relational patterns [119, 120, 121, 122, 123, 17, 19]. Building upon these advancements, we compute the relational embedding as $\mathbf{R}=f(\mathbf{X})\in\mathbb{R}^{V\times L\times D}$, which encapsulates the graph message at each time step. Here, $f$ is implemented as the Adaptive Graph Convolutional Network (AGCN) proposed in [122], allowing the skeleton graph topology to be learned adaptively. Subsequently, the relational embedding $\mathbf{R}$ is concatenated with the motion features and fed into the text-controlled selective SSM in Algorithm 2, to jointly capture the global temporal information of each joint alongside its topological context.
The overall architecture of the Text-Controlled Motion Mamba is depicted in Figure 7. Unlike the vanilla Mamba block, which adopts unidirectional causal modeling, the task of THMG demands the global context of the entire sequence. To address this issue, a bidirectional non-causal structure, as proposed in Vision Mamba [114], is employed. The input goes through a stack of TM-Mamba blocks, producing the output $\mathbf{Y}\in\mathbb{R}^{V\times L\times D}$. After mean pooling along the $V$ dimension, $\mathbf{Y}$ is subsequently forwarded to an MLP layer. This yields the frame activation score $s_{t}(t=1,2,\ldots,T)$ for each frame, indicating the likelihood of its inclusion in the retrieved temporal segments. The entire framework can be supervised using a simple cross-entropy loss.
$\mathcal{L}_{ce}=-\frac{1}{T}\sum_{t}^{T}(y_{t}\log{s_{t}}+(1-y_{t})\log{(1-s_{t})}).$ (5)
@@
V Experiments
V-A Implementation and Evaluation Details
The motion data pre-processing procedure adheres to the methodology outlined in BABEL [60]. The maximum length of the motion sequence is constrained to 2000. The same data split for training and evaluation as BABEL is adopted. CLIP [118] is adopted as text encoder to acquire the query embedding $q$ in Algorithm 2. To ensure fair comparison, we utilized the same text encoder throughout the experiments, namely clip-vit-base-patch32, with its parameters being jointly tuned together with the entire model during training. The feature dimension $D$ in Algorithm 2 for both the input motion sequence and text embedding is set to 256. The batch size and base learning rate are configured as $4$ and $5\times 10^{-4}$ respectively, with the learning rate for CLIP set to $5\times 10^{-5}$. The optimizer is an AdamW [124] with a weight decay of $1\times 10^{-4}$. The number of stacked Motion-Mamba blocks is empirically specified as $3$. Model training is conducted on a single Nvidia A40 GPU.
During the inference stage, a series of thresholds are used to obtain the predicted temporal segments, following [125]. Subsequently, non-maximum suppression is performed to remove overlapping segments. The evaluation of grounding performance is conducted using mean Average Precisions (mAPs) under different Intersection of Union (IoU) thresholds, namely $[0.1:0.7:0.1]$.
V-B Ablation Studies
@@
Text-Controlled Selection Mechanism. The text-controlled selection mechanism lies at the core of the TM-Mamba model. Removing this component directly from Algorithm 2 for ablation causes the model incapable of perceiving text query information. To enable comparison, the ablative models follow the practice of [58, 59], where text embeddings are concatenated with the sequence input. An MLP is then employed for feature fusion. As demonstrated in Table II, the text-controlled selection mechanism accounts for a substantial performance gain. This underscores the significance of dynamically regulating the propagation of information based on textual queries.
Unidirectional v.s. Bidirectional. The Vanilla Mamba model adopts a unidirectional approach, whereby the model can only access the sequence history while processing the current step. This unidirectional approach proves advantageous for tasks such as language modeling. However, in the context of the THMG task, the model requires a comprehensive understanding of the global context of the entire sequence, demanding a non-causal bidirectional structure. The results presented in Table II substantiate this claim, showing that the bidirectional model significantly outperforms its unidirectional counterpart.
@@
TABLE III: Performance comparisons to baseline methods on BABEL-Grounding dataset. The best results are in bold.
@@
S4D-LegS [96] 29.1 26.2 23.0 19.6 15.9 12.0 8.9 19.2
S4D-Lin [96] 30.1 26.9 23.8 20.5 16.6 12.8 9.5 20.0
2s-AGCN [122] 31.2 24.7 20.5 17.1 14.0 10.5 7.8 18.0
InfoGCN [120] 49.5 42.5 36.1 30.5 26.0 20.8 15.3 31.5
MotionCLIP [29] 41.8 39.2 36.1 32.7 29.3 24.9 19.9 32.0
MomentDETR [39] 51.1 46.0 39.1 32.9 26.8 20.5 13.6 32.9
EaTR [126] 53.4 48.5 43.6 37.3 31.0 23.5 16.2 36.2
STCAT [127] 47.1 44.2 40.7 37.1 33.3 28.9 23.6 36.4
@@
Figure 8: Comparison of memory consumption for TM-Mamba and its LSTM, classic Transformer and Flash-attention-2-based (FA2) Transformer counterparts under varying motion sequence lengths. The classic Transformer runs out of GPU memory when the sequence length reaches 1200.
TABLE IV: Performance comparison of TM-Mamba and its temporal Transformer counterpart under different maximum sequence length.
@@
Figure 9: Visualizations of predicted frame activation score. The solid red line denotes the predicted score of our full model, while the dashed blue and gray lines denote the model without text control and relational embeddings, respectively. The gray bars below illustrate the ground-truth temporal segments. Best viewed in color.
Figure 10: Visualizations of the grounding results of TM-Mamba compared to ablative models. ‘GT’ denotes ground-truth temporal segments corresponding to the text query. Our TM-Mamba demonstrates superior performance in motion grounding, in terms of the number of retrieved segments and the temporal boundaries of each segment. Best viewed in color.
V-C Comparison with Baselines
Given that the THMG task is novel and lacks existing works for comparison, several baseline methods are implemented for analysis. To begin, several prominent models are chosen from the video moment retrieval task: MomentDETR [39], EaTR [126] and STCAT [127]. However, these models rely on global temporal Transformers, which require quadratic computational memory and are impractical for handling the lengthy sequences in THMG task. As a remedy, their global temporal Transformers are substituted with efficient GRUs which require only linear memory. Also, two baselines based on powerful human motion backbones are implemented using spatial-temporal graph convolutions, namely 2s-AGCN [122] and InfoGCN [120]. Text embedding is incorporated through concatenation and MLP fusion to integrate query information for text-based grounding. We also implement a baseline where frame activation scores are computed using MotionCLIP [29], a motion-language model that aligns CLIP text embeddings with motion features through cosine similarity loss. Finally, we implemented several baselines based on recent SSM-based models [96] to demonstrate the advantages of TM-Mamba over other SSM methods. As illustrated in Table III, TM-Mamba outperforms the baseline models in terms of mAP at various IoU thresholds, thereby demonstrating its effectiveness.
V-D More Analysis